OpenStudy (miaaziz):

can somebody help me ? how to evaluate lim x approach to 0 .. (x-π)cot x ?

OpenStudy (sweetburger):

\[\lim_{x \rightarrow 0}(x-\pi )(\frac{ 1 }{ \tan(x) })\] plug in 0 \[\lim_{x \rightarrow 0}(0-\pi)(\frac{ 1 }{ \tan(0) })\] \[\lim_{x \rightarrow 0}\frac{ 1 }{ 0 }=?\]

OpenStudy (sweetburger):

*edit typo\[\lim_{x \rightarrow 0}\frac{ -\pi }{ 0 }\]

OpenStudy (miaaziz):

can i use the l'hospital's rule for this question ?

OpenStudy (sweetburger):

Actually I dont think you can as when you plug in x=0 it does not result in 0/0 or infinity/infinity

OpenStudy (irishboy123):

\[\lim_{x \rightarrow 0}(x-\pi) ~ \cot(x) \] \[= \lim_{x \rightarrow 0}(x-\pi)~~ \dfrac{ 1 }{ \tan(x) }\] \[= - \pi \lim\limits_{x \to 0} \dfrac{1}{ \tan(x) } \] no idea where it goes from here.

OpenStudy (irishboy123):

;-)

OpenStudy (miaaziz):

thanks